Zinc reacts with hydrochloric acid to form zinc chloride and hydrogen. What mass of zinc chloride could be produced if 25cm2 of hydrochloric acid of concentration 2.4 mol dm‐³ reacts with excess zinc


1
Expert's answer
2021-05-31T02:06:32-0400

The reaction could be expressed as:


"Zn+2HCl\rightarrow ZnCl_{2}+H_{2}"


The amount of HCl in moles is:


"\nu(HCl)=V(HCl)\cdot C(HCl)=0.025 dm^{3}\cdot2.4mol/dm^{3}=0.06 mol;"


The amount of ZnCl2 in moles is:


"\nu(ZnCl_{2})=\frac{1}{2}\cdot\nu(HCl)=0.03mol;"


The molar mass of ZnCl2 is:


"M(ZnCl_{2})=136g\/mol;"


Thus, the mass of ZnCl2 produced is:


"m(ZnCl_{2})=\nu(ZnCl_{2})\cdot M(ZnCl_{2})=4.08g."